Nolan walked to the store from his home at rate of 5 miles per hour. After spending one-half hour in the store, his friend gave him a ride home at the rate of 30 miles per hour. He arrived home 1 hour and 5 minutes (1 1/12 hours) after he left. How far is the store from Nolan's home?
Answer by stanbon(48514)

Nolan walked to the store from his home at rate of 5 miles per hour.
After spending one-half hour in the store, his friend gave him a ride home at the rate of 30 miles per hour.
He arrived home 1 hour and 5 minutes (1 1/12 hours) after he left.
How far is the store from Nolan's home?
-------------------------------------------
Walking DATA:
rate = 5 mph ; distance = x miles ; time = x/5 hrs
----------------------------------------------
Riding DATA:
rate = 30 mph ; distance = x miles ; time = x/30 hrs
---------------------------------------------------
Equation:
walk time + ride time = 13/12 hrs
x/5 + x/30 = 13/12
Multiply thru by 60 to get:
12x + 2x = 65
14x = 65
x = 65/14 miles is approximately 4.64
2x = 2(4.64) = 9.29 miles (distance from Nolan's house to the store)
===================================
Cheers,
Stan H.